# GazeNoter: Co-Piloted AR Note-Taking via Gaze Selection of LLM Suggestions to Match Users’ Intentions

### v1

@@ 5.2.3. Additional Discussions
For frustration, 6 participants (P1, P3, P4, P7, P10, P11) supposed that it was challenging to read text on the smartphone in (T) and (S) in the dynamic condition of walking. P4 further specified that “During walking, it was hard to maintain focus on the smartphone, read the texts and select keywords or sentences (S), not to mention organizing thoughts and then manually typing the notes (T).” Therefore, (G) significantly outperforms both (T) and (S) in frustration. P7, P10, and P11 mentioned that they gave up candidate sentence selection and directly recorded keywords as quick notes for less reading time. Consequently, the percentage of notes with complete sentences drops from 73% to 37% and the percentage of quick notes rises from 21% to 42% in (S). These disparities in (S) are much more pronounced compared to (T) and (G). This also leads to significantly better performance in reminding, quality, and inspiration in (G) compared to (S). For note refinement, no participants used this feature due to the interactive communication and dynamic condition in this scenario.
@@
In general, GazeNoter (G) significantly outperforms the other methods in all aspects. The requirement for occasional face-to-face discussions and the dynamic condition in walking meetings in this study result in obvious performance differences between GazeNoter on the smartphone (S) and the AR headset (G). Compared to smartphones, the absence of the need to look down, rapid switching across reality, and the large display are the primary benefits of AR headsets causing (G) to significantly outperform (S) in this study.
@@ 6. Limitations and Future Work
In general, GazeNoter was appreciated by participants in both studies. However, there are still some limitations. The accuracy of the built-in eye-tracking system in commercial AR headsets could occasionally disrupt the selection process, frustrating users. Using an advanced eye-tracking system in the future could improve not only frustration but also subtlety and social acceptance by reducing the time spent on diverted attention during selection. Another limitation is that real-world object or human tracking is not incorporated into the current system. Therefore, the current AR layout is not dynamically and automatically placed around the speaker, as in ParaGlassMenu (Cai et al., 2023). This will be achieved in the future work. To obtain beyond-context notes more precisely, the context and derivative keyword design in the current system could be further improved by using a fine-tuned model tailored to an individual’s knowledge, experiences, and even personality.
@@ 7. Conclusion
In this paper, we propose a real-time note-taking system, GazeNoter, by integrating a user-in-the-loop LLM system with gaze selection on an AR headset to generate notes that are both within-context and beyond-context, matching the users’ intentions. We evaluated GazeNoter’s effectiveness in two different scenarios: the static one where the user is stationary and listening to a speech, and the dynamic one where the user is walking and participating in a meeting. Our findings indicate that GazeNoter, through simple keyword and sentence selection processes on either smartphones or the AR headset, significantly outperforms both manual note-taking and automatic notes generated by the LLM in many metrics. Furthermore, when using the AR headset, GazeNoter demands less eye, head and hand movement, leading to less distraction and cognitive strain. More broadly, our work contributes to the ongoing efforts to integrate foundation models into XR applications.
Appendix A Prompts

### v2

@@ 5.2.3. Additional Discussions
Similar feedback and comments were mentioned for several factors as in the previous. We focus on discussing the differences caused by the mobile condition, interactive communication, and occasional face-to-face discussion in the walking meeting scenario. (G) significantly outperforms both (T) and (S) in the number of notes and time of a note. 6 participants (P2, P4, P9, P10, P11, P12) specified that the diverted attention caused by looking down at the phone in (T) and (S) during discussions led them to take fewer notes and spend less time on each note. This is also shown in Figure 17 and 18, demonstrating that the time users spent on each note composition was consistently longer in (G) compared to (S).
For social acceptance, P10 mentioned the necessity of seeking permission to take notes in (S) and (T) to avoid being perceived as impolite. P7 pointed out that they had to pause the discussion and sometimes even stop walking to take notes in (S) and (T), which hindered the ongoing conversation. As for (G), two users (P2 and P4) perceived the necessity to slightly move their eyes for interaction during face-to-face conversation diminishing social acceptance while others found it natural to glance away occasionally, enabling them to take notes. These also result in significantly higher social acceptance in (G) compared to (T) and (S), and underscore the influence of social settings on note-taking during occasional face-to-face discussions in walking meetings. Regarding subtlety and physical effort, P1, P4, P9, and P10 specified that constantly switching attention among the walking path, the phone, and the introducer with frequent head movement in (T) and (S) was physically demanding and less subtle. Thus, (G) significantly outperforms (T) and (S) in physical effort and subtlety. In addition, the longer duration for manual typing in (T) results in (T) being significantly less subtle than (S) in walking meetings.
@@
For frustration, 6 participants (P1, P3, P4, P7, P10, P11) supposed that it was challenging to read text on the smartphone in (T) and (S) in the mobile condition of walking. P4 further specified that “During walking, it was hard to maintain focus on the smartphone, read the texts and select keywords or sentences (S), not to mention organizing thoughts and then manually typing the notes (T).” Therefore, (G) significantly outperforms both (T) and (S) in frustration. P7, P10, and P11 mentioned that they gave up candidate sentence selection and directly recorded keywords as quick notes for less reading time. Consequently, the percentage of notes with complete sentences drops from 73% to 37% and the percentage of quick notes rises from 21% to 42% in (S). These disparities in (S) are much more pronounced compared to (T) and (G), as shown in the note type distribution of Figure 16 in the appendix. This also leads to significantly better performance in reminding, quality, and inspiration in (G) compared to (S). For (G), although the percentage of notes with complete sentences also drops from 78% to 63% and the percentage of quick notes rises from 18% to 26% (Figure 16), this difference is not as pronounced as in (S). P2, P7 and P9 mentioned that they would opt to record all selected keywords when they lacked the time to read the candidate sentences, such as in high-density information part of the speech-based activity. P1 and P5 also commented that they could obtain the desired sentences in notes when only selecting and recording keywords during such a situation. This demonstrates that participants can dynamically adapt their note-taking strategies based on different situations, balancing the trade-off between creating more complete and detailed longer notes and producing more concise and quicker notes. For note refinement, no participants used this feature due to the interactive communication and mobile condition in this scenario.
@@
In general, GazeNoter (G) significantly outperforms the other methods in all aspects. The requirement for occasional face-to-face discussions and the mobile condition in walking meetings in this study result in obvious performance differences between GazeNoter on the smartphone (S) and the AR headset (G). Compared to smartphones, the absence of the need to look down, rapid switching across reality, and the large display are the primary benefits of AR headsets causing (G) to significantly outperform (S) in this study.
@@ 6. Limitations and Future Work
In general, GazeNoter was appreciated by participants in both studies. However, there are still some limitations. The precision of the built-in eye-tracking system in commercial AR headsets could occasionally disrupt the selection process, frustrating users. Using an advanced eye-tracking system in the future could improve not only frustration but also subtlety and social acceptance by reducing the time spent on diverted attention during selection. Another limitation is that real-world object or human tracking is not incorporated into the current system. Therefore, the current AR layout is not dynamically and automatically placed around the speaker, as in ParaGlassMenu (Cai et al., 2023). This will be achieved in the future work. Furthermore, blurring and increasing transparency of items not selected or gazed at could reduce interference from AR content. However, the dynamic change and delays in these adjustments could interfere with users due to the rapid movement of the gaze.
While our system applies personal experiences and habits to customized keywords, it does not extend this to derivative keywords. To generate beyond-context notes more tailored to individual users, incorporating personal knowledge, experiences, and personality when creating derivative keywords could be beneficial. However, while tailored generation offers advantages, balancing user-specific outputs with out-of-the-box inspiration will be essential. The token limit of the LLM should be considered as well. Certainly, typing input could achieve the most precise notes. However, due to the real-time system design consideration (DC1), typing was not included in our design. The results of the studies also show that GazeNoter on both the smartphone and AR headset outperforms smartphone text entry. Nevertheless, using LLM to generate or auto-complete notes based on typed keywords is an advanced combination of the benefits of typing and our system, which generates notes (candidate sentences) based on selected keywords. Typing could be more precise but may cause more distraction. Therefore, comparing this and our system would be interesting in the future.
Although users can freely employ their note-taking strategies and two types of historical information are preserved, including previous context keywords in context keyword selection and transcripts in note refinement, to prevent missed content, we also envision that notes automatically generated by the LLM, as the baseline (A) in the studies, could be incorporated in the future. Although these notes may not always match users’ intentions and are limited to within-context, they can provide comprehensive contextual coverage to compensate for potential missed content. Thus, automatic note-taking (A) is complementary to our GazeNoter system.
@@ 7. Conclusion
In this paper, we propose a real-time note-taking system, GazeNoter, by integrating a user-in-the-loop LLM system with gaze selection on an AR headset to generate notes that are both within-context and beyond-context, matching the users’ intentions. We evaluated GazeNoter’s effectiveness in two different scenarios: the static one where the user is stationary and listening to a speech, and the mobile one where the user is walking and participating in a meeting. Our findings indicate that GazeNoter, through simple keyword and sentence selection processes on either smartphones or the AR headset, significantly outperforms both manual note-taking and automatic notes generated by the LLM in many metrics. Furthermore, when using the AR headset, GazeNoter demands less eye, head and hand movement, leading to less distraction and cognitive strain. More broadly, our work contributes to the ongoing efforts to integrate foundation models into XR applications.
Acknowledgements.
This research was supported in part by National Science and Technology Council of Taiwan (NSTC 111-2221-E-004 -008 -MY3, 113-2221-E-004 -007) and National Chengchi University.
Appendix A Additional objective result
Figure 12. Objective results of the note structures comparison between (S) and (G) in user study 1.
[…]
Objective results of the note structures comparison between (S) and (G) in user study 1.
Figure 13. The step distribution of note compositions of (S) and (G) in user study 1.
[…]
The step distribution of note compositions of (S) and (G) in user study 1.
Figure 14. The time distribution of note compositions of (S) and (G) in user study 1.
[…]
The time distribution of note compositions of (S) and (G) in user study 1.
Figure 15. Score from each within-context and beyond-context note in user study 1.
[…]
Score from each within-context and beyond-context note in user study 1.
Figure 16. Objective results of the note structures comparison between (S) and (G) in user study 2.
[…]
Objective results of the note structures comparison between (S) and (G) in user study 2.
Figure 17. The time distribution of note compositions of (S) and (G) in user study 2.
[…]
The time distribution of note compositions of (S) and (G) in user study 2.
Figure 18. The time distribution of note compositions of (S) and (G) in user study 2.
[…]
The time distribution of note compositions of (S) and (G) in user study 2.
Figure 19. Score from each within-context and beyond-context note in user study 2.
[…]
Score from each within-context and beyond-context note in user study 2.
Appendix B Prompts